Find the area of the region enclosed between the graphs of y=ex and y=e−x and the line x=ln2.
1
Expert's answer
2020-05-22T18:26:40-0400

The graphs of "y=e^x" and "y=e^{-x}" intersect if "x=0," therefore we should integrate "(e^x-e^{-x})" on the interval of [0, ln 2]:

"\int\limits_0^{\ln 2} (e^x-e^{-x})\,dx = (e^x + e^{-x})\Big|_{0}^{\ln 2} = (2+0.5) - (1+1) = 0.5."